OpenStudy (anonymous):

how do you solve, x/2+x/5=x/6-1/3

OpenStudy (eyust707):

You bring all the terms with x's to one side and the numerical term to the other side. Then you find the least common multiple of the denominators. You multiply each term by the appropriate factor to make the denominators all have the same LCM. Then you add the terms. Finally you multiply each side by the denominator so that x is isolated.

OpenStudy (anonymous):

can you explain further?

OpenStudy (theeric):

Well, you want x to be on only one side of the equation when you're finished, so that's actually a good place to start.

OpenStudy (theeric):

Given\[\frac{x}{2}+\frac{x}{5}=\frac{x}{6}-\frac{1}{3}\],

OpenStudy (anonymous):

X=5/8???

OpenStudy (anonymous):

i put all the x values on one side and i made the LCD 60 is that good so far?

OpenStudy (theeric):

Correct, GeekChic_! 60 is a good denominator! 30 also works if you'd like, but 60 will be fine too!

OpenStudy (anonymous):

yea thats what i got but shouldn't it be -5/8?

OpenStudy (anonymous):

yea lol i meant 30 as LCD

OpenStudy (theeric):

So you have\[\frac{30x}{60}+\frac{12x}{60}-\frac{10x}{60}=-\frac{1}{3}\]

OpenStudy (anonymous):

yup

OpenStudy (theeric):

Oh yeah, definately\[-\frac{5}{8}\]

OpenStudy (anonymous):

\[\frac{ 5x+2x }{ 10 }=\frac{ 3x-6 }{ 18 }\] then cross-multiply \[18(5x+2x)=10(5x+2x)\]

OpenStudy (anonymous):

x=-5/8

OpenStudy (anonymous):

alright thanks guys!

OpenStudy (theeric):

No problem! I'm glad you could work through the problem! Take care!